# Supersymmetric scale-separated AdS3 vacua of type IIB

### v1

@@ 2 A seven-dimensional solvmanifold, orbifolded and orientifolded
Table 1: The different orientifold planes after orientifolding the orbifold $\hat{\Gamma}$. The crosses indicate the directions parallel to the O5-plane, while the dashes correspond to the directions in which the O5-planes are localised, at the values $\{0,1/2\}$ and $\{1/4,3/4\}$, the latter being indicated with an asterisk.
Another example arises from introducing more shifts, for instance by quotienting with the $\check{\Gamma}=\mathbb{Z}_{2}^{3}$ group with the following group elements:
$\begin{split}&\check{\alpha}(x)=\left(-x^{1},-x^{2}+\frac{1}{2},x^{3},x^{4}+\frac{1}{2},-x^{5},-x^{6},x^{7}\right)\\ &\check{\beta}(x)=\left(x^{1},x^{2}+\frac{1}{2},-x^{3},-x^{4},-x^{5},-x^{6}+\frac{1}{2},x^{7}\right)\\ &\check{\gamma}(x)=\left(-x^{1}+\frac{1}{2},x^{2},-x^{3}+\frac{1}{2},x^{4},-x^{5}+\frac{1}{2},x^{6},-x^{7}\right)\,.\end{split}$ (8)
An orbifold like this appeared in ref. Andriolo:2018yrz but without the shifts in $\check{\gamma}$. Only the orbifold action $\check{\gamma}$ has fixed points, and again, Joyce’s criterion guarantees that this orbifold of the flat torus has singularities which can be blown up, of which there are only 4. Furthermore, only the orientifold involution $\sigma$ has fixed points, all orbifold images do not have any. Hence there is only one set of O5-planes. The orientifold involution does not further identify different orbifold singularities here either.
$x^{1}$ $x^{2}$ $x^{3}$ $x^{4}$ $x^{5}$ $x^{6}$ $x^{7}$
O5σ $-^{*}$$\times$ $-^{*}$$\times$ $-^{*}$$\times$ --
Table 2: The different orientifold planes after orientifolding the orbifold $\check{\Gamma}$. The crosses indicate the directions parallel to the O5-plane, while the dashes correspond to the directions in which the O5-planes are localised, at the values $\{0,1/2\}$ and $\{1/4,3/4\}$, the latter being indicated with an asterisk.
When the solvmanifold is quotiented by one of the orbifold groups $\Gamma$, $\hat{\Gamma}$ or $\check{\Gamma}$, the untwisted Betti numbers are reduced. Indeed, there are no one- or two-forms that are preserved under the orbifold, i.e. $b_{1}=b_{2}=0$ and there are only seven three-forms and seven four-forms that are preserved, i.e. $b_{3}=7$. A basis for them is given by
@@
where the short-hand notation $e^{a...b}=e^{a}\wedge...\wedge e^{b}$ is used. Note that all seven four-forms are closed under the exterior derivative by virtue of the Mauer-Cartan equations (3). The same is true for the first three four-forms, the other four are not closed. Consequentially, the seven-dimensional manifold admits a $G_{2}$-structure with a left-invariant three-form $\Phi$ and its associated four-form $\star\Phi$, which can be parametrised as follows:
$\begin{split}\Phi&=\;s_{127}\>e^{127}+s_{347}\>e^{347}+s_{567}\>e^{567}+s_{136}\>e^{136}+s_{235}\>e^{235}+s_{145}\>e^{145}-s_{246}\>e^{246}\\ \star\Phi&=\;s_{3456}\>e^{3456}+s_{1256}\>e^{1256}+s_{1234}\>e^{1234}+s_{2457}\>e^{2457}+s_{1467}\>e^{1467}\\ &\quad+s_{2357}\>e^{2367}-s_{1357}\>e^{1357}\,,\end{split}$ (10)
where $s_{ijk}=L_{i}L_{j}L_{k}$ and $s_{ijkl}=L_{i}L_{j}L_{k}L_{l}$. Again, the forms $\Phi$ and $\star\Phi$ are invariant under all orbifold actions $\Gamma$, $\hat{\Gamma}$ and $\check{\Gamma}$, and they are even under the orientifold involution $\sigma$ as well. The $G_{2}$ left-invariant three- and four-form do generally not need to be closed. Their non-closure can be described by the torsion classes, classified in different irreducible representations of the $G_{2}$ group:
@@
The Mauer-Cartan relations (3) imply that the 7d solvmanifolds admit a co-closed $G_{2}$-structure, but it does not necessarily have to be closed:
@@
However, by choosing the structure constants appropriately, i.e. $\omega_{\Sigma}=\omega_{a}+\omega_{b}+\omega_{c}=0$ and by taking $L_{1}=L_{2}$, $L_{3}=L_{4}$ and $L_{5}=L_{6}$, the solvmanifold has a closed three-form as well, which was already noticed in refs. Andriot:2015sia ; Manero2020315 . Nevertheless, co-closed $G_{2}$-structures play an important role in minimally supersymmetric type IIB compactifications to AdS3, as is discussed below.
@@ 3 Scale-separated solutions of type IIB with co-closed $G_{2}$-structures
This section uses the seven-dimensional solvmanifolds introduced above to construct AdS3 solutions of type IIB string theory. As they admit a co-closed $G_{2}$-structure, the tools of refs. Emelin:2021gzx ; VanHemelryck:2022ynr ; Passias:2020ubv can be used immediately. Compactifiying on such spaces preserves only 1/8 of the supersymmetries, and orientifolding breaks another half of them, resulting in minimal supersymmetry in three dimensions. The ten-dimensional metric can be written as a direct sum as follows:
@@
where $A$ is the warp factor, which is a function on the seven-dimensional space, governed by the metric $\mathrm{d}s_{7}^{2}$ which measures distances in string units. To find supersymmetric solutions, one can solve the killing spinor equations of type IIB string theory and impose the Bianchi identities. The killing spinor equations were rewritten in ref. Dibitetto:2018ftj in terms of bispinors for general backgrounds and applied in VanHemelryck:2022ynr ; Passias:2020ubv for $G_{2}$-structure backgrounds in type IIB. They require that the $G_{2}$-structure must be co-closed, a combination of the warping and the dilaton must be constant and that the NSNS three-form must vanish:
@@
$F_{1}=0$ (16)
$F_{3}=-g_{s}^{-1}\mathrm{e}^{-3A}\star\mathrm{d}\left(\mathrm{e}^{A}\Phi\right)+2\mu g_{s}^{-1}\mathrm{e}^{-3A}\Phi$ (17)
$F_{5}=0$ (18)
@@
The sum over $i$ runs over distinct O5-plane sets in different directions, while the sum over $j$ accounts for multiple O5-plane images parallel to those directions. It is useful to work in the smeared approximation for the O-planes, where the Dirac-delta forms are replaced by the unit volume forms of the cycles that the O5’s are orthogonal to, i.e. $\sum_{j}\delta(\Sigma_{\text{O5}_{\perp,i,j}})\to j_{\text{O5}_{i}}=\rho_{\mathrm{O5}}\text{vol}(\Sigma_{\text{O5}_{\perp,i}})$ where unit volume is understood here and $\rho_{\mathrm{O5}}$ is a constant charge density. In that case, the warping and dilaton profile are constant, and the solution for $F_{3}$ reduces to
@@
One concludes that, for any co-closed $G_{2}$-structure manifold, there exists a supersymmetric 3d compactification, provided that the manifold has the proper orientifold content such that the $F_{3}$-Bianchi identity is satisfied (in the smeared approximation).
As stated above, the different orbifolds $\hat{\Gamma}$ and $\check{\Gamma}$ of the seven-dimensional solvmanifold are examples of co-closed $G_{2}$-structures and can therefore be used for these type IIB compactifications. These are considered below.
Solutions on $\tilde{\mathbb{T}}^{7}/\hat{\Gamma}$
First, consider the space $\tilde{\mathbb{T}}^{7}/\hat{\Gamma}$, orbifolded by $\hat{\Gamma}$ of section 2 which, upon orientifolding, admits four sets of orientifold planes as in Table 1. It turns out to be useful to look at the situation where $L_{1}=L_{2}$, $L_{3}=L_{4}$ and $L_{5}=L_{6}$, which simplifies the setup but such that not all three-cycle volumes are the same. For this choice, the exterior derivative on $\Phi$ gives the simple expression:
It is important to note that when $\omega_{\Sigma}$ vanishes, the $G_{2}$-structure is both closed and co-closed. This was already noticed in refs. Andriot:2015sia ; Manero2020315 . In that case, $\mu$ also vanishes due to eq. (20) and the external spacetime is Minkowski instead of Anti-de Sitter. So for now, one can take $\omega_{\Sigma}\neq 0$. Even in this case, the Ricci curvature $R_{7}$ vanishes:
$R_{7}=\frac{21}{8}|W_{1}|^{2}-\frac{1}{2}|W_{27}|^{2}=0.$ (24)
Furthermore, the supersymmetric solutions above, eqs. (17)-(19) determine $F_{3}$ and $F_{7}$ completely in terms of the geometry:
Finally, $F_{3}$ is not closed, but due to the Mauer-Cartan equations (3), only the components including $\tilde{f}_{3}$ are responsible for that:
and since $\tilde{f}_{3}\omega_{\Sigma}>0$, the Bianchi identity of $F_{3}$ (21) should be solved by four sets of O5-planes. And indeed, as discussed in the previous sections, by orientifolding $\tilde{\mathbb{T}}^{7}/\hat{\Gamma}$, one gets, upon smearing, precisely the O5-planes required as in Table 1. Note also that the components of $F_{3}$ including the $f_{3,i}$’s are closed and do not participate in the Bianchi identity. This implies that all the $f_{3,i}$’s are unbounded flux quanta that can be chosen to be arbitrarily large.
Solutions on $\tilde{\mathbb{T}}^{7}/\check{\Gamma}$
Another interesting option is the scenario on $\tilde{\mathbb{T}}^{7}/\check{\Gamma}$, where the radii are also related in pairs by $L_{2}=\sqrt{2}L_{1}$, $L_{4}=\sqrt{2}L_{3}$ and $L_{6}=\sqrt{2}L_{5}$. Additionally, it is interesting to choose all structure constants to be the same in size, but where one of them has the opposite sign, which can be taken to be $\omega_{a}$ without loss of generality, i.e. $-\omega_{a}=\omega_{b}=\omega_{c}$. For this choice, the non-closure of $\Phi$ is captured by:
Note that there are some similarities but also key differences with the solution (25) on the previous orbifold. For example, the signs are different, especially the one for the 235-component, and the 246-component has an additional factor of 2. Nevertheless, the dependence on the flux parameters for the radii and string coupling is the same as what was considered for $\tilde{\mathbb{T}}^{7}/\hat{\Gamma}$. The RR three-form is not closed here either, and again, only the components including $\tilde{f}_{3}$ are responsible for that, only this time they generate just one component:
$\mathrm{d}F_{3}=-3\tilde{f}_{3}\,\omega_{a}\,\left(-e^{1357}\right)\,.$ (33)
To satisfy the Bianchi identity, only one set of (smeared) O5-planes is required. As indicated above, orientifolding $\tilde{\mathbb{T}}^{7}/\check{\Gamma}$ generates only one set of O5-planes along the proper directions, as in Table 2. Again, the $f_{3,i}$’s are not restricted by the tadpole and can be chosen to be arbitrarily large. Note that solutions on this space should be seen as the supersymmetric cousins of some of the solutions found in ref. Arboleya:2024vnp , which also include just one set of O5-planes and are non-supersymmetric. However, that paper uses another orbifold than the specific $\tilde{\mathbb{T}}^{7}/\check{\Gamma}$ and therefore might need additional D5-branes to cancel O5-charge along other directions.
Scale separation
Now it is time to investigate whether there is a separation of scales between the AdS scale and the volume of the three-cycles of the manifold. For this purpose, it is quintessential to make use of the fluxes that are unrestricted by tadpoles, which are parametrised here by the $f_{3,i}$’s and $f_{7}$. The flux parameter $\tilde{f}_{3}$ and structure constants $\omega_{a,b,c}$ are constrained by the $F_{3}$ Bianchi identity, as they have to be balanced against the O5-plane charge density. The latter is fixed, hence it cannot be tuned to infinity. Indeed, for both models, the Bianchi identity imposes that the product $\tilde{f}_{3}\omega_{\Sigma}$ for $\tilde{\mathbb{T}}^{7}/\hat{\Gamma}$ or $\tilde{f}_{3}\omega_{a}$ for $\tilde{\mathbb{T}}^{7}/\check{\Gamma}$ should be held fixed. However, all flux parameters are quantised, also $\tilde{f}_{3}$ and $\omega_{\Sigma}$ or $\omega_{a}$, which means that they should be held fixed separately: neither one of them can be made arbitrarily small at the expense of the other one becoming large. Another important remark is that the AdS scale is of similar size as the radius $L_{7}$, as
However, this is not a problem, as there are no cycles whose volumes are determined by $L_{7}$ alone. This is one of the reasons why the twisted seven-torus had to be orbifolded such that the one-forms, especially $e^{7}$, are projected out. The volumes of the generalised three-cycles are then set by $L_{i}^{2}L_{7}$ and $L_{1}L_{3}L_{5}$ for $i\in\{1,3,5\}$, which satisfy
where $f_{3,\alpha}$ is the smallest flux amongst the $f_{3,i}$’s. Hence scale separation can be achieved in the controlled regime, i.e. at small string coupling and large volume. Finally, it is also interesting to see how the 3d Planck mass scales with the fluxes when expressed in string units. In 3d, this is given by
$c=\frac{2}{3}M_{\text{Pl}}L_{\mathrm{AdS}}\sim f_{7}\;f_{3,1}\;f_{3,3}\;f_{3,5}\implies f_{3,\alpha}^{-2}(f_{3,1}\;f_{3,3}\;f_{3,5})^{2}\ll c\ll(f_{3,1}\;f_{3,3}\;f_{3,5})^{2}\,.$ (38)
This would result in a large number of degrees of freedom in the putative holographic dual theory. Additionally, there would be a parametrically large gap in the spectrum. Such conformal field theories with this feature have not been found yet.
Comments on using weak $G_{2}$-structures
So far, co-closed $G_{2}$-structures have been considered with non-vanishing torsion class $W_{27}$. This torsion class allows for non-positive scalar curvature, which requires O5-planes to solve Bianchi identities. However, it is also worthwhile to contemplate setups for which $W_{27}$ vanishes. In that case, only $W_{1}$ survives and the manifold is said to have a weak $G_{2}$-structure. Notice that $W_{1}$ is related to the inverse AdS radius and hence taking it to be vanishing would lead to Minkowski setups instead. So consider a type IIB background for which $W_{1}\neq 0$ and $W_{27}=0$. It follows from eq. (22) that
$\mathrm{d}F_{3}=g_{s}^{-1}\frac{W_{1}^{2}}{6}\star\Phi\,.$ (39)
Consequentially, the Bianchi identity should be solved by an appropriate amount of D5-branes wrapping three-cycles. The flux is not restricted in that case, as the number of D5-branes can be chosen at will. However, if one wants scale separation, then one would need an anisotropic manifold for which the torsion class $W_{1}$ and hence the scalar curvature radius decouple from the KK radius, estimated by the three- and four-cycle volume scales. Indeed, if the manifold is isotropic, then $\mu\sim W_{1}\sim 1/\mathcal{V}_{7}^{1/7}$, leading to no separation of scales. Furthermore, all homogeneous and smooth weak $G_{2}$-manifolds have been classified, and they do not allow for this property. Nevertheless, a weak $G_{2}$-structure manifold with these properties has recently been proposed in ref. VanHemelryck:2024bas , arising from a lift of a massless type IIA solution Cribiori:2021djm to M-theory, and could therefore lead to a hierarchy of scales in these 3d compactifications as well. However, exploring this scenario in full detail requires a more complete understanding of this space, which is left for future research.
4 Masses and conformal dimensions
@@
In refs. Conlon:2021cjk ; Apers:2022tfm ; Quirant:2022fpn , it was noticed that in the parametrically scale-separated AdS4 vacua of massive type IIA DeWolfe:2005uu ; Camara:2005dc , the masses of the scalars are such that their corresponding operators in the putative holographic field theory have integer conformal dimensions. The same feature arises in the scale-separated vacua of ref. Cribiori:2021djm , as they are morally T-dual to the vacua of refs.DeWolfe:2005uu ; Camara:2005dc , although on a different orbifold. In the three-dimensional vacua of ref. Farakos:2020phe , the situation is quite different, as it was shown in ref. Apers:2022zjx that the conformal dimensions are not integer there. Hence, it is worthwhile to investigate whether this feature also appears for the vacua constructed here.
@@
As was explained well in ref. Emelin:2021gzx , there are no axions in the type IIB setup here, either because homology does not allow for it (e.g. there is no $B$ or $C_{2}$ axion as there are no two-forms in the untwisted sector) or because the orientifold projects them out (e.g. $C_{4}$ is odd under an O5-involution but there are no odd four-cycles in the manifold). So the only scalar fields that enter the potential are the seven three-cylce volumes (or equivalently, the seven radii) and the dilaton. Since the radii $L_{k}$ defined above parametrise volumes in string units, it is convenient to trade the 10d dilaton $\phi$ for the 3d dilaton $\delta$. Additionally, it is useful here to trade the three-cycle volume scalars for the radii, such that the field space metric is diagonal:
@@
where the relation between both dilatons is that
@@
The supersymmetry conditions (15)-(19) are equivalent to the condition that $\partial_{i}P=0$ for all eight scalar fields. Again, it is more straightforward to calculate the mass matrix by trading $\phi$ for $\delta$. The masses are usually expressed in AdS units, therefore it is useful to observe that $V=\Lambda=-2L_{\mathrm{AdS}}^{-2}$ and hence

### v3

@@ 2 $\mathbb{Z}_{2}^{3}$ orbifolds of twisted tori and co-closed $G_{2}$-structures
[…]
Table 1: The different orientifold planes after orientifolding the orbifold $\hat{\Gamma}_{4}$. The crosses indicate the directions parallel to the O5-plane, while the dashes correspond to the directions in which the O5-planes are localised, at the values $\{0,1/2\}$ and $\{1/4,3/4\}$, the latter being indicated with an asterisk.
Another example arises from introducing more shifts, for instance, by choosing
$\hat{\Gamma}_{1}\;:\qquad a_{2}=a_{4}=b_{2}=b_{6}=c_{1}=c_{2}=c_{3}=1/2\,.$ (10)
An orbifold like this appeared in ref. Andriolo:2018yrz but without the shifts in $\hat{\gamma}$. Only the orbifold action $\hat{\gamma}$ has fixed points, and again, Joyce’s criterion guarantees that this orbifold of the ordinary torus has singularities which can be blown up, of which there are only 4. Furthermore, only the orientifold involution $\sigma_{\gamma}$ has fixed points, all orbifold images do not have any. Hence, there is only one set of O5-planes, summarised in Table 2. The orientifold involution does not further identify different orbifold singularities here either.
$\hat{\Gamma}_{1}$ $x^{1}$ $x^{2}$ $x^{3}$ $x^{4}$ $x^{5}$ $x^{6}$ $x^{7}$
[…]
Table 2: The different orientifold planes after orientifolding the orbifold $\hat{\Gamma}_{1}$. The crosses indicate the directions parallel to the O5-plane, while the dashes correspond to the directions in which the O5-planes are localised, at the values $\{0,1/2\}$.
The key idea is that by appropriately choosing the vectors $\vec{a}$, $\vec{b}$, and $\vec{c}$, along with a suitable orientifold involution $\sigma$, one has considerable flexibility in generating intersecting O5-planes along various directions. While the discussion so far has focused on orbifolds of the ordinary seven-torus $\mathbb{T}^{7}$, these constructions can often be extended to twisted tori $\tilde{\mathbb{T}}^{7}$. In such cases, however, the lattice by which one quotients does not need to be the square lattice, which is indeed the situation for nilmanifolds. When twisted tori are quotiented by any of the orbifold groups $\hat{\Gamma}$, the untwisted Betti numbers are reduced. Indeed, there are no one- or two-forms that are preserved under the orbifold, i.e. $b^{1}(\mathcal{M})=b^{2}(\mathcal{M})=0$, and there are only seven three-forms and seven four-forms that are preserved, i.e. $b^{3}(\mathcal{M})=7$. A basis for them is given by
@@
where the short-hand notation $e^{a...b}=e^{a}\wedge...\wedge e^{b}$ is used. Consequentially, the seven-dimensional manifold admits therefore a $G_{2}$-structure with a left-invariant three-form $\Phi$ and its associated four-form $\star\Phi$, which can be parametrised as follows:
[…]
where $s_{ijk}=L_{i}L_{j}L_{k}$ and $s_{ijkl}=L_{i}L_{j}L_{k}L_{l}$. Again, the forms $\Phi$ and $\star\Phi$ are invariant under all orbifold actions $\hat{\Gamma}$, and they are even under the orientifold involution $\sigma_{\hat{\gamma}}$ as well. The $G_{2}$ left-invariant three- and four-forms do generally not need to be closed. Their non-closure can be described by the torsion classes, classified in different irreducible representations of the $G_{2}$ group:
@@
Finally, note that all seven four-forms (11) are closed under the exterior derivative by the Maurer-Cartan equations (3) and the allowed structure constants (5). Consequentially, this means that the $\mathbb{Z}_{2}^{3}$ orbifolds lead to co-closed $G_{2}$-structures:
@@
Co-closed $G_{2}$-structures play an important role in minimally supersymmetric type IIB compactifications to AdS3, as is discussed below.
@@ 3 Scale-separated solutions of type IIB with co-closed $G_{2}$-structures
This section constructs AdS3 solutions of type IIB string theory, for which co-closed $G_{2}$-structures are essential. Solutions of this type have been studied before in refs. Passias:2020ubv ; Emelin:2021gzx and VanHemelryck:2022ynr , and the techniques used there apply directly here. Compactifications on these spaces preserve 1/8 of the original supersymmetry, and the inclusion of orientifold planes breaks half of the remaining supersymmetries, leading to minimal supersymmetry in three dimensions. The ten-dimensional metric can be written as a direct sum as follows:
@@
where $A$ is the warp factor, which is a function on the seven-dimensional space, governed by the metric $\mathrm{d}s_{7}^{2}$, which measures distances in string units. To find supersymmetric solutions, one can solve the Killing spinor equations of type IIB string theory and impose the Bianchi identities. The Killing spinor equations were rewritten in ref. Dibitetto:2018ftj in terms of bispinors for general backgrounds and applied in Passias:2020ubv ; VanHemelryck:2022ynr for $G_{2}$-structure backgrounds in type IIB. They require that the $G_{2}$-structure must be co-closed, a combination of the warping and the dilaton must be constant and that the NSNS three-form must vanish:
@@
$F_{1}=0\,$, (18)
$F_{3}=-g_{s}^{-1}\mathrm{e}^{-3A}\star\mathrm{d}\left(\mathrm{e}^{A}\Phi\right)+2\mu g_{s}^{-1}\mathrm{e}^{-3A}\Phi\,$, (19)
$F_{5}=0\,$, (20)
@@
The sum over $i$ runs over distinct O5-plane sets in different directions, while the sum over $j$ accounts for multiple O5-plane images parallel to those directions. In the rest of this work, the smeared approximation for the O-planes is imposed, where the Dirac-delta forms are replaced by the unit volume forms of the cycles that the O5’s are orthogonal to, i.e. $\sum_{j}\delta(\Sigma_{\text{O5}_{\perp,i,j}})\to j_{\text{O5}_{i}}=N_{\mathrm{O5}}\text{vol}(\Sigma_{\text{O5}_{\perp,i}})$ where unit volume is understood here and $N_{\mathrm{O5}}$ is the number of O5-planes. In that case, the warping and dilaton profile are constant, and the solution for $F_{3}$ reduces to
@@
One concludes that, for any co-closed $G_{2}$-structure manifold, there exists a supersymmetric 3d compactification, provided that the manifold has the proper orientifold content such that the $F_{3}$-Bianchi identity is satisfied (in the smeared approximation). As stated above, orbifolds of various seven-dimensional nilmanifolds and solvmanifolds are examples of co-closed $G_{2}$-structures and can therefore be used for these type IIB compactifications, which are discussed in the subsections below.
Masses and conformal dimensions
@@
Before examining specific solutions, the masses of the scalar fields and the conformal dimensions of their putative holographic dual operators are discussed here. In refs. Conlon:2021cjk ; Apers:2022tfm ; Quirant:2022fpn , it was noticed that in the parametrically scale-separated AdS4 vacua of massive type IIA DeWolfe:2005uu ; Camara:2005dc , the masses of the scalars are such that their corresponding operators in the putative holographic field theory have integer conformal dimensions. Although originally analysed for the $\mathbb{Z}_{3}\times\mathbb{Z}_{3}$ orbifold of DeWolfe:2005uu ; Camara:2005dc , it was shown in ref. Apers:2022zjx , that the same phenomenon occurs for other toroidal orbifolds, including $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$. Therefore, the same feature arises in the scale-separated vacua of ref. Cribiori:2021djm , as they are morally T-dual to the vacua of refs. DeWolfe:2005uu ; Camara:2005dc on the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ orbifold, as discussed in ref. Apers:2022zjx . In the three-dimensional vacua of ref. Farakos:2020phe , the situation is quite different, as it was shown in ref. Apers:2022zjx that the conformal dimensions are not integers there. Hence, it is worthwhile to investigate whether this feature also appears for the vacua constructed here.
@@
As was explained well in ref. Emelin:2021gzx , there are no axions in the type IIB setup here, either because homology does not allow for it (e.g. there is no $B$ or $C_{2}$ axion as there are no two-forms in the untwisted sector) or because the orientifold projects them out (e.g. $C_{4}$ is odd under an O5-involution but there are no odd four-cycles in the manifold). So the only scalar fields that enter the potential are the seven three-cycle volumes (or equivalently, the seven radii) and the dilaton. Since the radii $L_{k}$ defined above parametrise volumes in string units, it is convenient to trade the 10d dilaton $\phi$ for the 3d dilaton $\delta$. Additionally, it is useful here to trade the three-cycle volume scalars for the radii, such that the field space metric is diagonal:
@@
where the relation between the dilatons is that
@@
The supersymmetry conditions (17)-(21) are equivalent to the condition that $\partial_{i}P=0$ for all eight scalar fields. Again, it is more straightforward to calculate the mass matrix by trading $\phi$ for $\delta$. The masses are usually expressed in AdS units, therefore, it is useful to observe that $V=\Lambda=-2L_{\mathrm{AdS}}^{-2}$ and hence